If a seed is planted, it has a 85% chance of growing into a healthy plant. If 10 seeds are planted, what is the probability that exactly 3 don't grow?
Answer #1

Given that ,

p = 0.85

1 - p = 1 - 0.85 = 0.15

n =10

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 3) = (10 C 3) * 0.853 * (0.15)7

probability =0.0001